1.An investor Nx in bank M at the rate of 6%simple interest per annum and Ny in bank N at the rate of 8% simple interest per annum. If a total of #8,000,000.00 as interest from two banks after 4 years , calculate the (i) values of X and y; (ii) interest paid by the second bank


1
Expert's answer
2021-09-06T12:41:19-0400

Total amount invested 

"x+y=8,000,000.00"

Interest from bank "M = x×\dfrac{6}{100}×4=0.24x"


Interest from bank "N = y×\dfrac{8}{100}×4=0.32y"


The investor received a total of "A" as interest from the two banks after 4 years


"0.24x+0.32y=A"


"x+y=8000000""0.24x+0.32y=A"


"x=8000000-y""0.24(8000000-y)+0.32y=A"


"x=8000000-y""0.08y=A-1920000"


"x=8000000-y""y=12.5A-24000000"



"x=32000000-12.5A""y=12.5A-24000000"

(i)


"x=32000000-12.5A""y=12.5A-24000000"

(ii)


"0.32(12.5A-24000000)=4A-7680000"

Interest paid by the second bank is "N(4A-7680000)."


If "A=2,320,000.00," then


"x=3,000,000.00""y=5,000,000.00"


Interest paid by the second bank is "N1,600,000.00."